OpenStudy (anonymous):

PLEASE HELP! A geometry class has a total of 39 students. The number of males is 13 more than the number of females. How many males and how many females are in the class?

OpenStudy (anonymous):

first, we need to define any variables

OpenStudy (anonymous):

In a time crunch!!

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

let x represent the number of males let y represent the number of females

OpenStudy (anonymous):

what is the condition they have set for us? 'there are 13 more males than there are females' right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

hence x=13+y right? can you see this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so in total x+ y = 39 and we know x interns of y right? x=13+y so sub this into the equation (13+y)+y=39 13+2y=39 can you solve this now?

OpenStudy (anonymous):

y=13 x=26

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

i have another one

OpenStudy (anonymous):

no problem, fan and medal if you please

OpenStudy (anonymous):

The sum of two numbers is 49 and the difference is 9 . What are the numbers?

OpenStudy (anonymous):

we have two sets of equations. what does sum mean? and what does difference mean?

OpenStudy (anonymous):

sum implies + addition difference implies - subtraction, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if we have a number called x and a different number called y then x+y=49 and x-y=9

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can you solve these simultaneously now?

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

x=9+y right, just re-arrange the second equation. then sub this equation into the first equation. (9+y)+y=49 2y=40 y=20

OpenStudy (anonymous):

hence, x=29

OpenStudy (anonymous):

At a hockey game, a vender sold a combined total of 183 sodas and hot dogs. The number of sodas sold was two times the number of hot dogs sold. Find the number of sodas and the number of hot dogs sold.

OpenStudy (anonymous):

x+y=183 let x be soda y=hotdogs x=2y so 2y+y=183 3y=183 y=61, x=122

OpenStudy (anonymous):

the pattern is all similar, try this on your own now

OpenStudy (anonymous):

so number of hotdogs would be 61

OpenStudy (anonymous):

yep